solve each system of equations using the substitution method 5x+6y=-75x-3y=146
Answer by mananth(16949)

5x+6y=-75x-3y=146
You should separate the equations with a colon
5x+6y=-75 : x-3y=146
5x+6y=-75.............1
-3y=146 -1x
/-3
y=-48.67-0.33x
Plug the value of y in (1)
5x+6(-48.67+0.33 x)=-75
5x-292+2 x)=-75
5x+2x=-75+292
7x=217
/7
x=31
Plug the value of x in (1)
5x+6y=-75
5*31+6y =-75
155+6y=-75
6y=-230
/6
y=-38.33
CHECK
-3y=146 -1x
115=146 -31